suppose he price elasticity of demand for youe economics textbook is -1. If the publishr raises the price by 5% A. revenues will rise by 5% B. quantity demanded will rise 5% C. total revenues will not change D. revenues will fall
Answer #1

Answer

Option C

C. total revenues will not change D

the elasticity is -1 means it is unit elastic the change in price is equal to change in quantity.

TR=P*Q

after change

TR=1.05P * 0.95Q=0.9975*(PQ)=1*TR

the revenue will not change.